Consider the procedure of rolling a pair of dice 6 times and let x be the random variable consisting of the number of times the sum of the results is 7. The following table describes the probability distribution of x.

X

P(X)

0

0.334898

1

¿?

2

¿?

3

0.053584

4

0.008038

5

0.000643

6

0.000021

a) Find the missing probabilities

b) It would be unusual to roll a pair of dice six times and get at least three times the 7
Answer #1

bet xi number of times the sum is 7 xn Binomial (n=6, p. Pimf of X is P [x =] = 2 ( 2 ) pr (1 - Pjh-x , n = 0,3,2-06 Oow - GP ( x ) = 0.053584 +0.0080 + 0.0006 + 0. COOL ₃0.05 Since P(x2, 3 > 0.05 > not unsual to get at least three It is times the 7